# Traveling Waves Integrate Spatial Information into Spectral Representations

### v1

@@ 4 Results
Model Tetrominoes MNIST
@@
CNN 2-L 0.23 $\pm$ 0.00 0.16 $\pm$ 0.00 0.26 $\pm$ 0.00 0.27 $\pm$ 0.00 0.15 $\pm$ 0.01 0.24 $\pm$ 0.01
CNN 4-L 0.19 $\pm$ 0.00 0.26 $\pm$ 0.00 0.40 $\pm$ 0.00 0.22 $\pm$ 0.00 0.26 $\pm$ 0.00 0.38 $\pm$ 0.01
CNN 8-L 0.06 $\pm$ 0.01 0.70 $\pm$ 0.02 0.81 $\pm$ 0.02 0.13 $\pm$ 0.00 0.48 $\pm$ 0.01 0.62 $\pm$ 0.01
CNN 16-L 0.02 $\pm$ 0.01 0.92 $\pm$ 0.06 0.95 $\pm$ 0.03 0.04 $\pm$ 0.00 0.83 $\pm$ 0.01 0.89 $\pm$ 0.01
Conv-RNN-Last 0.65 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.69 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Conv-LSTM-Last 0.65 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.69 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Conv-RNN-FFT 0.20 $\pm$ 0.04 0.22 $\pm$ 0.10 0.33 $\pm$ 0.14 0.19 $\pm$ 0.10 0.41 $\pm$ 0.31 0.48 $\pm$ 0.30
Conv-LSTM-FFT 0.02 $\pm$ 0.01 0.96 $\pm$ 0.01 0.97 $\pm$ 0.01 0.10 $\pm$ 0.13 0.68 $\pm$ 0.40 0.71 $\pm$ 0.37
coRNN-FFT 0.09 $\pm$ 0.10 0.67 $\pm$ 0.37 0.72 $\pm$ 0.34 0.01 $\pm$ 0.00 0.95 $\pm$ 0.01 0.97 $\pm$ 0.01
@@
Table 1: Supervised segmentation performance of CNN baselines, Conv-RNNs, spectral representation methods, and U-Net, on Tetrominoes and MNIST. Models with loss $\leq 0.1$ are bolded.

### v2

@@ 4 Results
Model New Tetrominoes MNIST
@@
Baseline1 Flexible-2 0.23 $\pm$ 0.00 0.16 $\pm$ 0.00 0.26 $\pm$ 0.00 0.27 $\pm$ 0.00 0.15 $\pm$ 0.01 0.25 $\pm$ 0.00
Baseline1 Flexible-4 0.19 $\pm$ 0.00 0.26 $\pm$ 0.00 0.40 $\pm$ 0.00 0.22 $\pm$ 0.00 0.26 $\pm$ 0.01 0.39 $\pm$ 0.01
Baseline1 Flexible-8 0.06 $\pm$ 0.01 0.71 $\pm$ 0.00 0.82 $\pm$ 0.00 0.13 $\pm$ 0.00 0.50 $\pm$ 0.00 0.63 $\pm$ 0.00
Baseline1 Flexible-16 0.02 $\pm$ 0.01 0.92 $\pm$ 0.06 0.95 $\pm$ 0.03 0.04 $\pm$ 0.00 0.83 $\pm$ 0.01 0.89 $\pm$ 0.01
Baseline3 RNN-Last 0.65 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.70 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Baseline3 LSTM-Last 0.65 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.70 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Baseline3 FFT-RNN 0.10 $\pm$ 0.08 0.64 $\pm$ 0.31 0.72 $\pm$ 0.28 0.08 $\pm$ 0.06 0.73 $\pm$ 0.18 0.79 $\pm$ 0.16
Baseline3 FFT-LSTM 0.01 $\pm$ 0.00 0.97 $\pm$ 0.01 0.98 $\pm$ 0.00 0.04 $\pm$ 0.05 0.86 $\pm$ 0.16 0.89 $\pm$ 0.13
coRNN-FFT 0.01 $\pm$ 0.00 0.97 $\pm$ 0.00 0.98 $\pm$ 0.00 0.01 $\pm$ 0.00 0.95 $\pm$ 0.00 0.97 $\pm$ 0.00
@@
Table 1: Supervised segmentation performance of baseline models, and U-Net, on New Tetrominoes and MNIST. Models with loss $\leq 0.1$ are bolded.